OpenStudy (anonymous):

how do you find tan x/2 = sinx when 0≤x<2pi in radians? need step by step help!!!

jimthompson5910 (jim_thompson5910):

let z = x/2 So 2z = x x = 2z

jimthompson5910 (jim_thompson5910):

which means we get tan(z) = sin(2z)

OpenStudy (anonymous):

tan x/2 = sin2x?

jimthompson5910 (jim_thompson5910):

Then we can use the identities tan(A) = sin(A)/cos(A) and sin(2A) = 2*sin(A)cos(A) -------------------------- tan(z) = sin(2z) sin(z)/cos(z) = 2*sin(z)cos(z) sin(z)/cos(z) - 2*sin(z)cos(z) = 0 sin(z)( 1/cos(z) - 2*cos(z)) = 0 sin(z) = 0 or 1/cos(z) - 2*cos(z) = 0

jimthompson5910 (jim_thompson5910):

Solve those equations for z. Then use those solutions to find x.

jimthompson5910 (jim_thompson5910):

Do you see how I'm getting all this?

OpenStudy (anonymous):

z=x/2 right?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

how do you get sin(z) / cos(z) to sin(z)( 1/cos(z) ?

jimthompson5910 (jim_thompson5910):

I factored out sin(z)

OpenStudy (anonymous):

could you show me the ssteps:(??? i dont know how!!

OpenStudy (anonymous):

wait i will draw it. could you please check!?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

im so lost....

jimthompson5910 (jim_thompson5910):

one sec

jimthompson5910 (jim_thompson5910):

what you drew isn't valid (i don't think)

OpenStudy (anonymous):

but you cant factor sin(z)( 1/cos(z) - 2*cos(z)) = 0 because then the denominator becomes sinzcosz....

OpenStudy (anonymous):

so far we have this!

jimthompson5910 (jim_thompson5910):

\[\Large \tan(z) = \sin(2z)\] \[\Large \frac{\sin(z)}{\cos(z)} = 2\sin(z)\cos(z)\] \[\Large \frac{\sin(z)}{\cos(z)} - 2\sin(z)\cos(z) = 0\] \[\Large \sin(z)\left(\frac{1}{\cos(z)} - 2\cos(z)\right) = 0\] With me so far ?

OpenStudy (anonymous):

is it like this??

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

We can go further... \[\Large \tan(z) = \sin(2z)\] \[\Large \frac{\sin(z)}{\cos(z)} = 2\sin(z)\cos(z)\] \[\Large \frac{\sin(z)}{\cos(z)} - 2\sin(z)\cos(z) = 0\] \[\Large \sin(z)\left(\frac{1}{\cos(z)} - 2\cos(z)\right) = 0\] \[\Large \sin(z)\left(\frac{1}{\cos(z)} - \frac{2\cos^2(z)}{\cos(z)}\right) = 0\] \[\Large \sin(z)\left( \frac{1-2\cos^2(z)}{\cos(z)}\right) = 0\] \[\Large \sin(z)\left( \frac{-1(2\cos^2(z)-1)}{\cos(z)}\right) = 0\] \[\Large \sin(z)\left( \frac{-cos(2z)}{\cos(z)}\right) = 0\] \[\Large \sin(z)=0 \ \text{or} \ \frac{-cos(2z)}{\cos(z)} = 0\] \[\Large \sin(z)=0 \ \text{or} \ -cos(2z) = 0\] \[\Large \sin(z)=0 \ \text{or} \ cos(2z) = 0\]

OpenStudy (anonymous):

i didnt get this step @jim_thompson5910 !!

jimthompson5910 (jim_thompson5910):

it should be 2cos^2(z) - 1 that becomes cos(2z) See this page for trig identities http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (anonymous):

what....im still lost >< im so sorry!

jimthompson5910 (jim_thompson5910):

one of the trig identities on that page is \[\Large \cos(2u) = 2\cos^2(u)-1\] do you see it

jimthompson5910 (jim_thompson5910):

you have to scroll a bit

OpenStudy (anonymous):

ahh okay!!

jimthompson5910 (jim_thompson5910):

so do you see how i got all that?

OpenStudy (anonymous):

im kind of lost on -cos2(z) / cos(z) = 0 how did it become just -cos2(z) = 0 ?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

I multiplied both sides by the denominator cos(z)

jimthompson5910 (jim_thompson5910):

Basically, we have something like A/B = 0 Multiply both sides by B to get A = 0*B A = 0

OpenStudy (anonymous):

oh yesah ! okay!!

jimthompson5910 (jim_thompson5910):

so your goal is to solve sin(z) = 0 and cos(2z) = 0

OpenStudy (anonymous):

yes so since z= x/2 ill just plug in sinx/2 = 0

OpenStudy (anonymous):

then itll be sin x = 2?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

sin(x/2) = 0 does NOT become sin(x) = 2

jimthompson5910 (jim_thompson5910):

sin(z) = 0 sin(x/2) = 0 x/2 = arcsin(0) x/2 = 0 or x/2 = pi x = 0 or x = 2pi

jimthompson5910 (jim_thompson5910):

cos(2z) = 0 cos(2(x/2)) = 0 cos(x) = 0 x = arccos(0) x = pi/2 or x = 3pi/2

OpenStudy (anonymous):

does arcsin = \[\sin^-1\]

jimthompson5910 (jim_thompson5910):

yes, arcsine and inverse sine are the same thing

jimthompson5910 (jim_thompson5910):

now we have the following possible solutions x = 0, x = pi/2, x = pi, x = 2pi

jimthompson5910 (jim_thompson5910):

but when x = pi, tan(x/2) = tan(pi/2) which is undefined So x = pi is NOT a solution

jimthompson5910 (jim_thompson5910):

also, x = 2pi isn't a solution either because it's not in the interval 0≤x<2pi

OpenStudy (anonymous):

wait...x/2 = 0 or x/2 = pi i dont get this....

jimthompson5910 (jim_thompson5910):

So only \[\Large x = \frac{\pi}{2}\] and \[\Large x =\frac{3\pi}{2}\] are solutions

jimthompson5910 (jim_thompson5910):

oh i had sin(x/2) = 0 and I took the arcsine of both sides to get x/2 = arcsine(0) then I evaluated the arcsine of 0 to get 0 or pi Note: sin(0) = 0 and sin(pi) = 0

OpenStudy (anonymous):

i got it!

jimthompson5910 (jim_thompson5910):

ok great

OpenStudy (anonymous):

@jim_thompson5910 im sorry again but could you please make it clear about the very beginning. you put z = x/2 so we plug in x/2 for all the (z) right ?

jimthompson5910 (jim_thompson5910):

yes because we let z = x/2 at the top we replaced all copies of "z" with "x/2"

jimthompson5910 (jim_thompson5910):

since z = x/2, we can multiply both sides by 2 to get x = 2z. So we can replace all copies of 'x' with '2z'

jimthompson5910 (jim_thompson5910):

Then we simplify/solve to get sin(z) = 0 or cos(2z) = 0

jimthompson5910 (jim_thompson5910):

from there, we can go ahead and replace 'z' with 'x/2' (since z = x/2) and 2z with x (since x = 2z) to get sin(x/2) = 0 or cos(x) = 0

jimthompson5910 (jim_thompson5910):

Does that make sense?

OpenStudy (anonymous):

yes thank you!!!

jimthompson5910 (jim_thompson5910):

alright great